# Conditional and Modal Reasoning in Large Language Models

### v3

@@ 4.2 Inconsistency and overgeneralization
Most interestingly, we tested a number of inference patterns which involve the interaction of modals with conditionals or disjunction. This is an especially interesting domain, since work in philosophy and logic has shown that substituting a modal sentence for a non-modal one can change a pattern from being apparently valid to apparently invalid.
@@
These cases are of special interest to test on LLMs since human subjects can immediately recognize that while DS, MT, and MP are valid for Boolean sentences, they apparently fail for modal/conditional substitution instances of these patterns like DSmu, MTmu, and CMP. That is, humans do not overgeneralize from the simple Boolean case to the general case. It is thus interesting to explore whether LLMs are human-like in this respect or rather overgeneralize from simple to complex cases—especially since these complex cases have only been discussed in a relatively small number of philosophy and logic papers and are presumably somewhat rare in naturalistic settings and hence presumably not very frequent in training data.
We found that, indeed, many LLMs do overgeneralize: they do not agree with human judgments about the invalidity of MTmu, DSmu, and CMP. Intriguingly, some models exhibit human-like judgments in rejecting MTmi, but at the same time they accept MTmu (Figure 2). This is logically inconsistent, since those models also accept that ‘might not’ is logically equivalent to ‘not must’ (MiN and NMu), [^3] In fact, the inconsistenty arises just using MiN. For given MiN, the premises of MTmi entail the premises of MTmu, so one cannot reject the former and accept the latter. and MTmi and MTmu differ only by substituting these equivalent phrases. This intriguing pattern suggests that LLMs may indeed overgeneralize from the validity of MT to judging MTmu to be valid, while recognizing that MTmi (which is not syntactically an instance of MT) is invalid.
Figure 2: Zero-shot responses for MTmu (above) and MTmi (below) show inconsistency for many models.
@@
Figure 4: Percentage of responses that were jointly consistent when we asked leading models about DSmu, MiN, and DSmi at the same time, in some order. The large standard deviations (error bars) show strong sensitivity to question orders, which is highly undesirable.
@@
Figure 6: Correlations of our evaluation results (zero-shot) vs. LMSYS Elo ratings, MMLU scores, and GSM8k scores. The correlations are 0.74, 0.77, and 0.72, respectively. All p-values are less than 0.01.
We found similar patterns for DS, as shown in Figure 3, where most models accepted DSmu but rejected DSmi. Indeed, even the GPT-4 models we tested exhibit logically inconsistent behavior in this area, accepting DSmu and MiN while rejecting DSmi. This inconsistency was displayed for both zero-shot and chain-of-thought prompts. We wondered whether this incoherence could be due to the fact that the inconsistent triad of responses was given in response to different prompts in different context windows, so we probed more deeply, asking the leading models about each of the three inference patterns within the same context window, but their responses remained logically inconsistent much of the time; see Figure 4 for a summary.
@@
Due to space constraints, graphs showing results for all other inferences are included in Appendix C.
@@ 4.3 Relationship to some popular benchmarks
To position our logical reasoning tasks with respect to the broad landscape of LLM evaluations, we compare the models’ performance on our benchmark (non-controversial inferences) to that of Chatbot Arena (general assistance, Chiang et al., 2024), MMLU (domain knowledge, Hendrycks et al., 2020), and GSM8K (math reasoning, Cobbe et al., 2021), all of which are popular benchmarks to assess LLM capabilities. We show that our results are highly correlated with the results of each of the three in Figure 6. The Arena Elo ratings come from LMSYS directly. [^4] https://chat.lmsys.org The MMLU scores are combined from those given by LMSYS and Open LLM Leaderboard, [^5] https://huggingface.co/spaces/open-llm-leaderboard/open_llm_leaderboard both using a 5-shot setup. The GSM8K scores are obtained from Zhang et al. (2024). The high correlations are not surprising, but it supports the hypothesis that logical reasoning abilities are related to and predictive of not only mathematical reasoning abilities but also domain-general capabilities. It would be interesting in future work to investigate the causal connections: whether improving logical reasoning also improves some general reasoning/assistance abilities.
@@ 5 Discussion
On the one hand, our results mirror what we have learned in the field over the past few years: larger models are likely to perform better at reasoning, and chain-of-thought prompting often but does not always help. On the other hand, we have identified inconsistent and counterintuitive reasoning behaviors from even the best models with and without chain-of-thought. The inference patterns that give rise to those behaviors reflect state-of-the-art research in philosophy and logic, which by nature means they are less present in the training and fine-tuning data of LLMs (though most likely not absent). This is suggestive of sources from which we can acquire more out-of-distribution evaluation data to test LLMs, and it illustrates that LLMs’ judgments may not be reliable when they encounter novel inference patterns, even if those are natural and intuitive for humans. Lastly, we note that a neurosymbolic, semantic-parsing approach to logical reasoning, as in Olausson et al. 2023, will not automatically handle the tasks we present here. To our knowledge, theorem provers do not natively implement the nuanced meaning representations we discussed in this paper, and in some cases it is still controversial what the correct semantics for certain modal operators are. Thus, working towards an automated natural language reasoning system that sensibly solves our tasks is a challenge regardless of the models and approaches.
@@ 6 Conclusion
We have drawn on philosophical, logical, and linguistic analyses of inference to explore the extent to which LLMs are accurate logical reasoners about conditionals and modals. We hope this will be the start of a new strand of research on LLMs. There are many natural follow-ups. First, it would be interesting to compare the behavior of LLMs with experimental human subjects on all the inferences we tested. We have reported expert human judgments from the literature, but the judgments of experimental subjects might exhibit mistakes interestingly similar to or different from those we find in LLMs. Second, there are many more inference patterns to explore, involving modals, conditionals, as well as many other parts of the human logical apparatus, like generalized quantifiers, attitude predicates, and degree constructions. Third, modal and conditional reasoning is also naturally connected to other rich topics concerning reasoning, such as probabilistic reasoning, causal models, and mental simulations. The language of modals and conditionals could provide another perspective for studying such reasoning in LLMs and humans. More generally, we hope that there will be more work studying LLMs using the insights of philosophy, linguistics, and logic into the fundamental building blocks of human reasoning and meaning.
@@ 7 Limitations
In this paper, we have focused on studying to what extent LLMs understand logical inferences with conditionals and modals. The conditional connective and ‘must’ and ‘might’ operators are the target, while we do not systematically study other important logical operators such as ‘and’, ‘or’, or quantifiers such as ‘some’ and ‘all’. Even in the modal domain, operators like ‘probably’ and ‘certainly’ also deserve careful analysis. In particular, the interactions of many of these different kinds of operators could lead to interesting inference patterns, which would be worthy of future study. Additionally, we employ simple syntactical constructions to create the evaluation data. While this is natural for testing logical inference, the models’ performance on an inference pattern may not generalize when it is made of more complex phrases. Lastly, our dataset is in English, and the models’ logical inference abilities may differ by language as well. We hope future work could study the multilingual aspect of logical inference as well.
@@ Acknowledgments
For helpful feedback, we thank Dominic Hughes, James Kirkpatrick, Hailey Schoelkopf, and Shane Steinert-Threlkeld.
@@ Appendix A Language models used
In this section, we cite all the LLMs used to conduct our experiments: the GPT-4 model family OpenAI (2023); the Claude 3 model family; [^6] https://www.anthropic.com/news/claude-3-family Gemini 1.5 Pro and Flash Reid et al. (2024); Llama 3 Instrcut 70B and 8B; [^7] https://ai.meta.com/blog/meta-llama-3/ Mixtral 8x7B Jiang et al. (2024) and Mistral 7B Jiang et al. (2023); Llama 2 Chat 7B, 13B, and 70B Touvron et al. (2023); Code Llama 7B, 13B, and 34B Roziere et al. (2023); Yi Chat 34B Young et al. (2024); Phi 2.888https://www.microsoft.com/en-us/research/blog/phi-2-the-surprising-power-of-small-language-models/
@@ Appendix B Prompts used in the experiments
We use the following prompts to run the experiments reported in this work. The prompts are the same for all models, except that for some small models (Phi-2, Llama 2 7B, and Code Llama 7B), we had to modify the chain-of-thought prompt with extra reminders to answer only after first thinking step by step. Each prompt is shown with a concrete example of an inference pattern.
@@ B.2 Few-shot
From Ann went to the store’, does it follow that Ann went to the store and Bob went to the beach’? No, it does not follow.
From Ann went to the store’, does it follow that Ann went to the store or Bob went to the beach’? Yes, it does follow.
From Ann went to the store and Bob went to the beach’, does it follow that Ann went to the store’? Yes, it does follow.
From Ann went to the store or Bob went to the beach’, does it follow that Ann went to the store’? No, it does not follow.
@@ Appendix C Additional results
C.1 Performance summaries
Figure 7: Summary of performance on the simple logical inference patterns under different conditions and temperature 0.
Figure 8: Summary of performance on the simple logical inference patterns under zero-shot and few-shot conditions with temperature 1.
C.2 Individual inferences
In this Appendix, we display the zero-shot performance of all LLMs on the inferences shown in Table 2.3 when operating at both temperature 0 and temperature 1.
C.2.1 Disjunctive Syllogism (DS)
@@ C.2.11 Contraposition (CT)
C.2.12 DS with ‘must’ (DSmu)
@@
C.2.13 DS with ‘might’ (DSmi)
@@
C.2.14 MT with ‘must’ (MTmu)
@@
C.2.15 MT with ‘might’ (MTmi)
@@
C.2.16 Complex Modus Ponens (CMP)

### v4

@@ 4.2 Inconsistency and overgeneralization
We tested a number of inference patterns that involve the interaction of modals with conditionals or disjunction. This is an especially interesting domain, since work in philosophy and logic has shown that substituting a modal sentence for a non-modal one can change a pattern from being apparently valid to apparently invalid.
@@
These cases are of special interest to test on LLMs since human subjects can immediately recognize that while DS, MT, and MP are valid for Boolean sentences, they apparently fail for modal/conditional substitution instances of these patterns like DSmu, MTmu, and CMP. That is, humans do not overgeneralize from the simple (Boolean) case to the general case. It is thus interesting to explore whether LLMs are human-like in this respect or rather overgeneralize from simple to complex cases—especially since these complex cases have only been discussed in a relatively small number of philosophy and logic papers and are presumably somewhat rare in naturalistic settings and hence presumably not very frequent in training data.
We found that, indeed, many LLMs do overgeneralize: they do not agree with human judgments about the invalidity of MTmu, DSmu, and CMP. Intriguingly, some models exhibit human-like judgments in rejecting MTmi, but at the same time they accept MTmu (Figure 2). This is logically inconsistent, since those models also accept that ‘might not’ is logically equivalent to ‘not must’ (MiN and NMu),  [^3] In fact, the inconsistenty arises just using MiN. For given MiN, the premises of MTmi entail the premises of MTmu, so one cannot reject the former and accept the latter. and MTmi and MTmu differ only by substituting these equivalent phrases. This intriguing pattern suggests that LLMs may indeed overgeneralize from the validity of MT to judging MTmu to be valid, while recognizing that MTmi (which is not syntactically an instance of MT) is invalid. [^4] Our examples were designed to elicit an epistemic reading of the modals in question. However, this is immaterial to our logical points, since even if the target LLM instead accessed a deontic reading, the judgments we report would remain jointly inconsistent, as long as the LLM’s interpretation does not change across different instances of the modal in question.
Figure 2: Zero-shot responses for MTmu (above) and MTmi (below) show inconsistency for many models. All error bars, including in subsequent figures, represent 95% confidence intervals.
@@
Figure 4: Percentage of responses that were jointly consistent when we asked leading models about DSmu, MiN, and DSmi in the same context window, in one of the six possible orders. Each dot represents such an order. The results show strong sensitivity to question order, which is highly undesirable.
@@
Figure 6: Correlations of our evaluation results (zero-shot) vs. LMSYS Elo ratings, MMLU scores, and GSM8k scores. The correlations are 0.81, 0.85, and 0.75, respectively. All p-values are less than 0.01.
We found similar patterns for DS, as shown in Figure 3, where many models accepted DSmu but rejected DSmi. Indeed, even the GPT-4 models we tested exhibit logically inconsistent behavior in this area, accepting DSmu and MiN while rejecting DSmi. This inconsistency was displayed for both zero-shot and chain-of-thought prompts. We wondered whether this incoherence could be due to the fact that the inconsistent triad of responses was given in response to different prompts in different context windows, so we probed more deeply, asking the leading models about each of the three inference patterns within the same context window, but their responses remained logically inconsistent much of the time; see Figure 4 for a summary.
@@
4.3 Modal fallacies and free choice
We also found intriguing patterns involving LLMs’ behavior with purely modal inferences, shown in Appendix C.1.17-C.1.20. While, as noted, the models were generally able to correctly reason about duality (inferring ‘not must’ from ‘might not’, and vice versa), they systematically made basic errors in modal reasoning in other cases. Strikingly, almost all models, including those that performed best overall, accepted both MuDistOr and MiAg as valid, despite these being clearly invalid. (For MuDistOr, asssume the keys must be upstairs or downstairs; it doesn’t follow that either they must be upstairs or they must be downstairs. After all, they might not be upstairs, and they might not be downstairs. For MiAg, assume the keys might be upstairs and might be downstairs; it doesn’t follow that they might be both upstairs and downstairs.) The situation with WSFC and NSFC is also interesting. There is controversy in the semantics literature about the status of these patterns, which strike many speakers as valid but are not valid on standard modal semantics Kamp (1973). Intriguingly, many models accepted NSFC but rejected WSFC. In fact, this conforms to one position in the literature which maintains that, indeed, NSFC but not WSFC is valid Simons (2005); Meyer and Sauerland (2017); Fusco (2019). However, this threatens inconsistency, since most models also accept the (standardly valid) equivalence of $\lozenge(p\vee q)$ with $\lozenge p\vee\lozenge q$ (see the GitHub repository).
@@ 4.4 Relationship to some popular benchmarks
To position our logical reasoning tasks with respect to the broad landscape of LLM evaluations, we compare the models’ performance on our benchmark (uncontroversial inferences) to that of Chatbot Arena (general assistance, Chiang et al., 2024), MMLU (domain knowledge, Hendrycks et al., 2020), and GSM8K (math reasoning, Cobbe et al., 2021), all of which are popular benchmarks to assess LLM capabilities. We show that our results are highly correlated with the results of each of the three in Figure 6. The Arena Elo ratings come from LMSYS directly. [^5] https://chat.lmsys.org The MMLU and GSM8k scores are obtained from the HELM leaderboard Liang et al. (2022). [^6] https://crfm.stanford.edu/helm The high correlations support the hypothesis that logical reasoning abilities are related to and predictive of not only mathematical reasoning abilities but also domain-general capabilities. It would be interesting to investigate the causal connections: whether improving logical reasoning also improves some general reasoning abilities.
@@ 5 Discussion
On the one hand, our results mirror what we have learned in the field over the past few years: larger models are likely to perform better at reasoning, and chain-of-thought prompting often but does not always help. On the other hand, we have identified inconsistent and counterintuitive reasoning behaviors from even the best models with and without chain-of-thought. The inference patterns that give rise to those behaviors reflect state-of-the-art research in philosophy and logic, which by their nature means they are less present in the training and fine-tuning data of LLMs (though most likely not absent). This is suggestive of sources from which we can acquire more out-of-distribution evaluation data to test LLMs, and it illustrates that LLMs’ judgments may not be reliable when they encounter novel inference patterns, even if those are natural and intuitive for humans. Lastly, we note that a neurosymbolic, semantic-parsing approach to logical reasoning, as in Olausson et al. 2023, will not automatically handle the tasks we present here. To our knowledge, theorem provers do not natively implement the nuanced meaning representations we discussed in this paper, and in some cases it is still controversial what the correct semantics for certain modal operators are. Thus, working towards an automated natural language reasoning system that sensibly solves our tasks is a challenge regardless of the models and approaches.
@@ 6 Conclusion
We have drawn on philosophical, logical, and linguistic analyses of inference to explore the extent to which current LLMs are accurate logical reasoners about conditionals and modals. We hope this will be the start of a new strand of research on LLMs. There are many natural follow-ups. First, it would be interesting to compare the behavior of LLMs with experimental human subjects on all the inferences we tested. We have reported expert human judgments from the literature, but the judgments of experimental subjects might exhibit mistakes interestingly similar to or different from those we find in LLMs. Second, there are many more inference patterns to explore, involving modals, conditionals, as well as many other logical constructions, like quantifiers, attitude predicates, and degree constructions. Third, modal and conditional reasoning is also connected to probabilistic reasoning, causal models, and mental simulations, and hence could provide another perspective for studying these in LLMs and humans. More generally, we hope that there will be more work studying LLMs using philosophical, linguistic, and logical insights into the building blocks of reasoning and meaning.
@@ 7 Limitations
In this paper, we have focused on studying to what extent LLMs understand logical inferences with conditionals and modals. The conditional connective and ‘must’ and ‘might’ operators are the target, while we do not systematically study other important logical operators such as ‘and’ and ‘or’, or quantifiers such as ‘some’ and ‘all’. Even in the modal domain, operators like ‘probably’ and ‘certainly’ also deserve careful analysis. In particular, the interactions of many of these different kinds of operators could lead to interesting inference patterns, which would be worthy of future study. Additionally, we employ simple syntactical constructions to create the evaluation data. While this is natural for testing logical inference, the models’ performance on an inference pattern may not generalize when it is made of more complex phrases. Lastly, our dataset is in English, and the models’ logical inference abilities may differ by language. We hope future work could study the multilingual aspect of logical inference as well.
@@ Acknowledgments
For helpful feedback, we thank Dominic Hughes, James Kirkpatrick, Roger Levy, Hailey Schoelkopf, Shane Steinert-Threlkeld, and the reviewers for EMNLP 2024.
@@ Appendix A Language models used
In this section, we cite all the LLMs used to conduct our experiments: the GPT-4 model family OpenAI (2023); the GPT-3.5 model family Brown et al. (2020); [^7] https://openai.com/index/chatgpt the Claude 3 model family [^8] https://www.anthropic.com/news/claude-3-family and Claude 3.5 Sonnet; [^9] https://www.anthropic.com/news/claude-3-5-sonnet Gemini 1.5 Pro and Flash Reid et al. (2024); Gemma 2 27B Gemma Team (2024); the Llama 3 model family Dubey et al. (2024); Mistral Large 2; [^10] https://mistral.ai/news/mistral-large-2407 Mixtral 8x7B Jiang et al. (2024) and Mistral 7B Jiang et al. (2023); Llama 2 Chat 7B and 13B Touvron et al. (2023); Code Llama 7B, 13B, and 34B Roziere et al. (2023).
@@ Appendix B Prompts used in the experiments
We use the following prompts to run the experiments reported in this work. The prompts are the same for all models, except that for some small models (e.g., Llama 2 7B and Code Llama 7B), we had to modify the chain-of-thought prompt with extra reminders to answer only after first thinking step by step. Each prompt is shown with a concrete example of an inference pattern.
@@ B.2 Few-shot
From Ann went to the store’, can we infer that Ann went to the store and Bob went to the beach’? Correct answer: No.
From Ann went to the store’, can we infer that Ann went to the store or Bob went to the beach’? Correct answer: Yes.
From Ann went to the store and Bob went to the beach’, can we infer that Ann went to the store’? Correct answer: Yes.
From Ann went to the store or Bob went to the beach’, can we infer that Ann went to the store’? Correct answer: No.
@@ Appendix C Additional results
C.1 Individual inferences
In this Appendix, we display the zero-shot performance of the LLMs on the inferences shown in Table 2.3 when operating at both temperature 0 and temperature 1.
C.1.1 Disjunctive Syllogism (DS)
@@ C.1.11 Contraposition (CT)
C.1.12 DS with ‘must’ (DSmu)
@@
C.1.13 DS with ‘might’ (DSmi)
@@
C.1.14 MT with ‘must’ (MTmu)
@@
C.1.15 MT with ‘might’ (MTmi)
@@
C.1.16 Complex Modus Ponens (CMP)
@@
C.1.17 ‘Must’ distribution over ‘or’ (MuDistOr)
[Uncaptioned image][Uncaptioned image]
C.1.18 ‘Might’ agglomeration (MiAg)
[Uncaptioned image][Uncaptioned image]
C.1.19 Narrow-scope free choice (NSFC)
[Uncaptioned image][Uncaptioned image]
C.1.20 Wide-scope free choice (WSFC)
[Uncaptioned image][Uncaptioned image]
C.2 Performance summaries
Figure 7: Summary of performance on the uncontroversial logical inference patterns under different conditions and temperature 0.
Figure 8: Summary of performance on the uncontroversial logical inference patterns under zero-shot and few-shot conditions with temperature 1.